OpenStudy (anonymous):

HELP MEEE!!!! Solve the system of equations by determinants for z only: 2x+y=3z=-2 x-4y+z=24 -3x-y+4z=0

OpenStudy (anonymous):

Ive found the original D and the \[D_z\]

OpenStudy (anonymous):

I just dont understand the dividing part

OpenStudy (anonymous):

@study100 can you help?

OpenStudy (anonymous):

2x +y =3z =2? is it 2x + y + 3z = 2

OpenStudy (anonymous):

That's how it is on the paper, I also thought this was a typo, made by the professor. @study100

OpenStudy (anonymous):

@study100 what do you think it might be? cause it could be 2x+y-3z=2 OR 2x+y+3z=2

OpenStudy (anonymous):

Yes that might be the case but I think it's +3z, since the + and = are on the same key

OpenStudy (anonymous):

yea lets just go with that

OpenStudy (anonymous):

Now I have to form the quotient for D_z and D

OpenStudy (anonymous):

but i'm not sure how to get that from this

OpenStudy (anonymous):

2 1 3 D: 1 4 1 -3 -1 4

OpenStudy (anonymous):

2 1 -2 D_z: 1 4 24 -3 -1 0